# The Atacama Cosmology Telescope: Release of A databaSe of millimeTeR ObservatioNs of Asteroids Using acT (ASTRONAUT)

### v1

@@ Abstract
1 Introduction
Studies of asteroids in the solar system have increased our understanding of early planetary formation processes as diverse asteroid spectra have indicated a dynamic solar system history (DeMeo & Carry, 2014). Studying the composition of asteroid sub-surfaces (commonly referred to as the asteroid regolith) can further increase our understanding of physical processes in the nascent solar system. Asteroid measurements at millimeter wavelengths provide a way to study regolith composition since these measurements are sourced from below the asteroid surface (Ulich & Conklin, 1976; Johnston et al., 1982; Viikinkoski et al., 2015). Different asteroid sub-surface compositions can lead to varying flux measurements as a function of wavelength. Comparing these measurements to models in the infrared (IR) have suggested a sub-millimeter (sub-mm) deficit (e.g., results from Müller & Barnes (2007); Orlowski-Scherer et al. (2024); Chichura et al. (2022)).
Changes in asteroid composition can signify the presence of temperature gradients or changes in emissivity as a function of wavelength, although more work is needed to better explain which process is responsible. Millimeter observations can therefore shed light on the formation of the asteroid belt in the solar system. However, there is a lack of millimeter flux data from asteroids compared to other wavelengths. Most asteroid studies have been performed in the optical and IR regimes, requiring large observing resources for targeted searches (Müller & Barnes, 2007; Chamberlain et al., 2007). Millimeter observations are also limited by the apparent size of asteroids, as well as their approximately black body spectral radiance profile. Additionally, asteroids are quite faint at these wavelengths compared to the optical/infrared wavelengths where most asteroid science is done. This is both due to their spectral slope and the diffraction limit imposing a $10^{3}-10^{6}$ times worse beam dilution factor. Millimeter observations are therefore relatively costly, and have until recently been limited to targeted observations of a few asteroids.
@@ 1 Introduction
To increase access to these data, we present A databaSe of millimeTeR ObservatioNs of Asteroids Using acT (ASTRONAUT) hosted by Amazon Web Services, Inc. (AWS). This database is in the form of a public Amazon Simple Storage Service (S3) bucket. This bucket contains normalized thermal emission flux measurements, associated error bars, weighting factors based on asteroid geometry, and observation times. To aid in the S3 bucket user implementation, we also include a Jupyter notebook, which is hosted on the ASTRONAUT GitHub repository.  [^1] https://github.com/ACTCollaboration/ASTRONAUT This notebook details methods to pull the flux data and create light curves.
In this paper, we present data products from an analysis of 170 asteroids extracted from observations made by ACT in frequency bands centered near 98, 150, and 228 GHz, denoted f090, f150, and f220, respectively (Orlowski-Scherer et al., 2024). ACT can probe into the regolith with depths of a few millimeters to a few centimeters, although these distances are influenced by different material properties. The asteroids presented in this study contain values at the $5$ sigma signal-to-noise level or higher in at least one observing band when stacking over all observations.
This paper is structured as follows. In Section 2, we provide an overview of the ACT data and briefly describe the depth-1 maps used to study asteroids. In Section 3, we summarize the asteroid data products from the S3 bucket and describe the bucket itself. We also discuss light curves, the main purpose for this container. In Section 4, we present a Jupyter Notebook which describes how to use the bucket and create an example light curve. Finally, in Section 5, we summarize this work.
2 The Atacama Cosmology Telescope
@@
2.1 Depth-1 Maps
@@
2.2 Data Processing
We measure the asteroid fluxes from the depth-1 maps following Orlowski-Scherer et al. (2024). The depth-1 maps are matched filtered for point sources, which cuts out atmospheric noise scales as well as scales smaller than the beam size. This matched-filter maximizes the signal-to-noise (S/N) for point sources on the maps, which includes asteroids. Orlowski-Scherer et al. (2024) then create small stamps of depth-1 maps using the JPL Horizons API. [^2] https://ssd.jpl.nasa.gov/horizons/ This API was implemented to determine the location of asteroids on the depth-1 maps using the ACT observation times. Using the Horizons ephemerides, it was possible to interpolate the orbit of asteroids provided observing times. Once these asteroids are centered, measuring the flux corresponds to evaluating the matched-filter flux map value at the location of the asteroid on a single depth-1 map. The flux error is calculated from a flux uncertainty map. These flux values can be used to create light curves, as described in Section 3.2. This database therefore contains the flux values returned from the center of depth-1 stamps after a normalization procedure, described below. This also includes the flux error, as well as the weighting values used to derive the normalized fluxes. The time corresponds to the observation time of the asteroid, up to the array crossing time (approximately $4$ minutes), which is returned in Unix time units. Unix time is a unit of measurement corresponding to the number of non-leap seconds that have passed since 00:00:00 UTC on January 1, 1970 (Stevens & Rago, 2013). We refer readers to Orlowski-Scherer et al. (2024) for more details.
Figure 1: Example plot of the depth-1 maps discussed in Section 2.1 which are used to generate light curves discussed in Section 3.2. Light curve data shown here are for (4) Vesta on array pa5 at $150$ GHz. The color bar corresponds to flux intensity. Note that this data release does not contain data maps such as those shown here.
3 Data Products
@@
Table 1: Table of ACT array/frequency combinations and array sensitivities (Naess in prep.) supported by this database.
ACT Array Frequency (GHz) Sensitivity ($\mu K\sqrt{s}$)
pa4 $150,220$ $19.1,42.6$
pa5 $90,150$ $12.6,15.1$
pa6 $90,150$ $11.4,13.6$
3.1 Amazon Simple Storage Service
@@ 3.2 Light Curves
We discourage users from comparing asteroids based on their time-averaged signal. Specifically, averaging across different asteroid depth-1 maps will not yield interpretable results due to differences in regolith composition. To average the data, we recommend placing the asteroids into specific classes, which is discussed in Orlowski-Scherer et al. (2024).
3.3 Available Asteroids
@@
4 Jupyter Notebook
In addition to launching this database, we also create a Jupyter Notebook to aid users. This Notebook guides users through requesting data from the public bucket (described in Section 3), reading in the data file(s), and generating an example light curve.
@@ 5 Conclusion
Future applications of these data include creating light curves across observatories, such as ACT, SPT, and the Atacama Large Millimeter/submillimeter Array, as well as calculating phase curves to better understand the variation of millimeter flux as a function of asteroid rotation. These data could also be used to calibrate millimeter instruments such as those mentioned above.
6 Acknowledgments
This work was supported by the U.S. National Science Foundation through awards AST-1440226, AST0965625 and AST-0408698 for the ACT project, as well as awards PHY-1214379 and PHY-0855887. The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. ACT operates in the Parque Astronómico Atacama in northern Chile under the auspices of the Comisión Nacional de Investigación Científica y Tecnológica de Chile (CONICYT), now La Agencia Nacional de Investigación y Desarrollo (ANID). Colleagues at AstroNorte and RadioSky provide logistical support and keep operations in Chile running smoothly.Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund–Research Excellence, and the University of Toronto.
@@ 6 Acknowledgments
We thank Bez Thomas for their technical support.

### v2

@@ Abstract
I Introduction
Studies of asteroids in the solar system have increased our understanding of early planetary formation processes as diverse asteroid spectra have indicated a dynamic solar system history (DeMeo and Carry, 2014). Asteroid measurements at millimeter wavelengths provide a way to study regolith composition since these measurements are sourced from the top of the asteroid subsurface (Ulich and Conklin, 1976; Johnston et al., 1982; Viikinkoski et al., 2015). Different asteroid sub-surface compositions can lead to varying flux measurements as a function of wavelength. Comparing these measurements to models in the infrared (IR) have suggested a sub-millimeter (sub-mm) deficit (e.g., results from Müller and Barnes (2007); Orlowski-Scherer et al. (2024); Chichura et al. (2022)).
Changes in asteroid composition can signify the presence of temperature gradients or changes in emissivity as a function of wavelength, although more work is needed to better explain which process is responsible. Millimeter observations can therefore constrain regolith properties of asteroids. However, there is a lack of millimeter flux data from asteroids compared to other wavelengths. Asteroid observations in the high-frequency radio regime have been around for decades (Conklin et al., 1977; Johnston et al., 1982; Hunter et al., 2015; Shepard et al., 2021; Ostro et al., 1985), and most asteroid studies have been performed in the optical and IR regimes, requiring large observing resources for targeted searches (Müller and Barnes, 2007; Chamberlain et al., 2007). Millimeter observations are also limited by the apparent size of asteroids, as well as their approximately black body spectral radiance profile. Additionally, asteroids are quite faint at these wavelengths compared to the optical/infrared wavelengths where most asteroid science is done. This is both due to their spectral slope and the diffraction limit imposing a $10^{3}-10^{6}$ times worse beam dilution factor. Millimeter observations are therefore relatively costly, and have until recently been limited to targeted observations of a few asteroids.
@@ I Introduction
To increase access to these data, we present A databaSe of millimeTeR ObservatioNs of Asteroids Using acT (ASTRONAUT) hosted by Amazon Web Services, Inc. (AWS). This database is in the form of a public Amazon Simple Storage Service (S3) bucket. This bucket contains normalized thermal emission flux measurements, associated error bars, weighting factors based on asteroid geometry, and observation times. To aid in the S3 bucket user implementation, we also include a Jupyter notebook, which is hosted on the ASTRONAUT GitHub repository.  [^1] https://github.com/ACTCollaboration/ASTRONAUT This notebook details methods to pull the flux data and create light curves. We note that the use of the term “light-curve” in this paper is non-standard in the asteroid community. However, this paper is intended for a wider audience, so we keep this usage as it is standard terminology outside of this field.
In this paper, we present data products from an analysis of 170 asteroids extracted from observations made by ACT in frequency bands centered near 98, 150, and 228 GHz, denoted f090, f150, and f220, respectively (Orlowski-Scherer et al., 2024). ACT can probe the regolith to depths of millimeters to a few centimeters, although these penetration scales into the composite media are directly influenced by wavelength, material properties, composition, and the detailed geometry of the particles in the media (Sihvola, 2008; Bohren and Huffman, 2004). The asteroids presented in this study contain values at the $5$ sigma signal-to-noise level or higher in at least one observing band when stacking over all observations.
This paper is structured as follows. In Section II, we provide an overview of the ACT data and briefly describe the depth-1 maps used to study asteroids. In Section III, we summarize the asteroid data products from the S3 bucket and describe the bucket itself. We also discuss light curves, the main purpose for this container. In Section IV, we present a Jupyter Notebook which describes how to use the bucket and create an example light curve. Finally, in Section V, we summarize this work.
II The Atacama Cosmology Telescope
@@
Table 1: Table of ACT array/frequency combinations, array sensitivities (Naess et al., 2025), and corresponding beam areas supported by this database. The areas are given at the ACT beam’s full width at half maximum (Duivenvoorden, in preparation).
ACT Array Frequency (GHz) Sensitivity ($\mu K\sqrt{s}$) Beam Area (arcmin2)
pa4 $150,220$ $19.1,42.6$ $2.45$, $1.27$
pa5 $90,150$ $12.6,15.1$ $5.19$, $2.42$
pa6 $90,150$ $11.4,13.6$ $5.19$, $2.48$
II.1 Depth-1 Maps
@@
II.2 Data Processing
We measure the asteroid fluxes from the depth-1 maps following Orlowski-Scherer et al. (2024). The depth-1 maps are matched filtered for point sources, which cuts out atmospheric noise scales as well as scales smaller than the beam size. This matched-filter maximizes the signal-to-noise (S/N) for point sources on the maps, which includes asteroids. Orlowski-Scherer et al. (2024) then create small stamps of depth-1 maps using the JPL Horizons API. [^2] https://ssd.jpl.nasa.gov/horizons/ This API was implemented to determine the location of asteroids on the depth-1 maps using the ACT observation times. Using the Horizons ephemerides, it was possible to interpolate the orbit of asteroids provided observing times. Once these asteroids are centered, measuring the flux corresponds to evaluating the matched-filter flux map value at the location of the asteroid on a single depth-1 map. The flux error is calculated from a flux uncertainty map. These flux values can be used to create light curves, as described in Section III.2. This database therefore contains the flux values returned from the center of depth-1 stamps after a normalization procedure, described below. This also includes the flux error, as well as the weighting values used to derive the normalized fluxes. The time corresponds to the observation time of the asteroid, up to the array crossing time (approximately $4$ minutes), which is returned in Unix time units. Unix time is a unit of measurement corresponding to the number of non-leap seconds that have passed since 00:00:00 UTC on January 1, 1970 (Stevens and Rago, 2013). We refer readers to Orlowski-Scherer et al. (2024) for more details.
Figure 1: Example plot of the depth-1 maps discussed in Section II.1 which are used to generate light curves discussed in Section III.2. Light curve data shown here are for (4) Vesta on array pa5 at $150$ GHz. The color bar corresponds to flux intensity. Note that this data release does not contain data maps such as those shown here.
III Data Products
@@
III.1 Amazon Simple Storage Service
@@ III.2 Light Curves
III.3 Available Asteroids
@@
IV Jupyter Notebook
In addition to launching this database, we also create a Jupyter Notebook to aid users. This Notebook guides users through requesting data from the public bucket (described in Section III), reading in the data file(s), and generating an example light curve.
@@ V Conclusion
Future applications of these data include creating light curves across observatories, such as ACT, SPT, and the Atacama Large Millimeter/submillimeter Array, as well as calculating phase curves to better understand the variation of millimeter flux as a function of asteroid rotation.
VI Acknowledgments
This work was supported by the U.S. National Science Foundation through awards AST-1440226, AST0965625 and AST-0408698 for the ACT project, as well as awards PHY-1214379 and PHY-0855887. The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. ACT operates in the Parque Astronómico Atacama in northern Chile under the auspices of the Comisión Nacional de Investigación Científica y Tecnológica de Chile (CONICYT), now La Agencia Nacional de Investigación y Desarrollo (ANID). Colleagues at AstroNorte and RadioSky provide logistical support and keep operations in Chile running smoothly. Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund-–Research Excellence, and the University of Toronto.
@@ VI Acknowledgments
We thank Bez Thomas and Kailai Wang for their technical support.